# Nonlinearly Preconditioned Gradient Methods under Generalized Smoothness

### v1

@@ Appendix B Missing proofs of Section 2
Note that the function $\nabla\phi^{*}\circ(\bar{L}^{-1}\nabla f)$ is locally Lipschitz, since $\nabla\phi^{*}$ is globally Lipschitz and $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$. Then, we can invoke the generalized mean value theorem in its summation form [facchinei2003finite, Proposition 7.1.16]: for two points $x,\bar{x}\in\mathbb{R}^{n}$, there exist $n$ points $z_{i}\in(x,\bar{x})$ and $n$ scalars $\alpha_{i}\geq 0$ summing to unity such that
where $V_{i}\in\partial_{C}(\nabla\phi^{*}\ \circ\bar{L}^{-1}\nabla f)(z_{i})$. Now, in light of [clarke1990optimization, p. 75], $\partial_{C}(\nabla\phi^{*}\circ\bar{L}^{-1}\nabla f)(x)v\subseteq\operatorname{con}\{\partial_{C}(\nabla\phi^{*})(\bar{L}^{-1}\nabla f(x))\bar{L}^{-1}\nabla^{2}f(x)v\}$ for any $v\in\mathbb{R}^{n}$. Therefore, any $V_{i}(x-\bar{x})$ can be written as $\bar{L}^{-1}\sum_{j=1}^{d}\beta_{j}H_{j}\nabla^{2}f(z_{i})(x-\bar{x})$, for $d>0$, with $H_{j}\in\partial_{C}(\nabla\phi^{*})(\bar{L}^{-1}\nabla f(z_{i}))$ and $\beta_{j}\geq 0$ summing to unity.
Taking an inner product, we have that
$\langle\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x))-\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\bar{x})),x-\bar{x}\rangle$ $=\bar{L}^{-1}\sum_{i=1}^{n}\alpha_{i}\langle x-\bar{x},\sum_{j=1}^{d}\beta_{j}H_{j}\nabla^{2}f(z_{i})(x-\bar{x})\rangle$
$=\bar{L}^{-1}\sum_{i=1}^{n}\alpha_{i}\sum_{j=1}^{d}\beta_{j}\langle x-\bar{x},H_{j}\nabla^{2}f(z_{i})(x-\bar{x})\rangle$
$\leq\bar{L}^{-1}\sum_{i=1}^{n}\alpha_{i}\bar{L}L\|x-\bar{x}\|^{2}$
$=L\|x-\bar{x}\|^{2},$
@@
which is the claimed result.
For the reverse implication, we have that under Assumption 1.2, $T_{{L^{-1}},{\bar{L}^{-1}}}(x)=x-L^{-1}\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x))$ is a differentiable mapping. Then, since it is also monotone, in light of [RoWe98, Proposition 12.3], its Jacobian $T_{{L^{-1}},{\bar{L}^{-1}}}(x)=I-L^{-1}\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))\bar{L}^{-1}\nabla^{2}f(x)$, which is not necessarily symmetric, is positive-semidefinite, implying Definition 2.4.
B.5 Proof of Corollary 2.7
Since $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$ is $(L_{0},L_{1})$-smooth, we have that $\lambda_{\max}(\nabla^{2}f(x))\leq L_{0}+L_{1}\|\nabla f(x)\|$. For $\nabla\phi^{*}(y)=\frac{y}{1+\|y\|}$, $L=L_{1}$ and $\bar{L}=L_{0}/L_{1}$, we have that
$T_{{\delta L^{-1}},{\bar{L}^{-1}}}(x)$ $=x-\tfrac{\delta}{L_{1}}\frac{\nabla f(x)\tfrac{L_{1}}{L_{0}}}{1+\tfrac{L_{1}\|\nabla f(x)\|}{L_{0}}}$
$=x-\frac{\delta\nabla f(x)}{L_{0}+L_{1}\|\nabla f(x)\|}.$
Moreover, we have that $L\bar{L}[\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))]^{-1}=L_{0}[\nabla^{2}\phi^{*}(L_{1}L_{0}^{-1}\nabla f(x))]^{-1}\succeq L_{0}(1+L_{1}L_{0}^{-1}\|\nabla f(x)\|)I=(L_{0}+L_{1}\|\nabla f(x)\|)I$. Therefore, $\lambda_{\max}(\nabla^{2}f(x))\leq L\bar{L}\lambda_{\min}([\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))]^{-1})$ and the result follows from Lemma 2.5.
@@
Note that (4) is equivalent to $\bar{x}\in\operatorname*{arg\,min}_{x\in\mathbb{R}^{n}}g(x):=\bar{L}(L^{-1}\star\phi)(x-\bar{y})-f(x)$ for all $\bar{x}\in\mathbb{R}^{n}$, where $\bar{y}=\bar{x}-L^{-1}\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\bar{x}))$. In light of the modern version of Fermat’s theorem [RoWe98, Theorem 10.1], for $\tilde{x}$ to be a local minimizer of $g$, the following inclusion should hold: $0\in\widehat{\partial}g(\tilde{x})$. Through [RoWe98, Exercise 10.10] this implies that
@@
or that $L(\tilde{x}-\bar{y})=\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\tilde{x}))$ or $\tilde{x}-L^{-1}\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\tilde{x}))=\bar{y}$. This means that $T_{{L^{-1}},{\bar{L}^{-1}}}(\tilde{x})=T_{{L^{-1}},{\bar{L}^{-1}}}(\bar{x})$ and since $T_{{L^{-1}},{\bar{L}^{-1}}}$ is strictly monotone, the only possible minimizer of $g$ is $\bar{x}$.
@@
Case 1: $\operatorname{dom}\phi$ is bounded. In light of [RoWe98, p. 91], $g$ is a coercive function and as it is also closed, it attains its minimum.
@@
$\leq\bar{L}r^{-1}\alpha\phi\left(\frac{r}{\alpha}(x-\bar{y})\right)+\bar{L}r^{-1}(1-\alpha)\phi\left(\frac{r}{1-\alpha}\bar{y}\right)$
$-\frac{\mu}{2}\bar{L}r^{-1}\alpha(1-\alpha)\left\|\frac{r}{\alpha}(x-\bar{y})-\frac{r}{1-\alpha}\bar{y}\right\|^{2},$
@@
Note now that $\psi$ is a closed and strongly convex function and as such it has bounded level-sets. Due to the inequality, the level-sets of $g$ are contained in those of $\psi$ and thus $g$ also has bounded level-sets. Since moreover $g$ is closed, it attains its minimum.
@@ Appendix C Missing proofs of Section 3
$f(x^{k+1})\leq f(x^{k})-(1-|1-\alpha|)\bar{L}L^{-1}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k}))).$
@@
$\sum_{k=0}^{K}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k})))\leq\frac{L}{\bar{L}(1-\beta)}(f(x^{0})-f(x^{K+1}))\leq\frac{L}{\bar{L}(1-\beta)}(f(x^{0})-f^{\star}),$
@@
$(K+1)\min_{0\leq k\leq K}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k})))\leq\frac{L}{\bar{L}(1-\beta)}(f(x^{0})-f^{\star}).$ (18)
@@
$\min_{0\leq k\leq K}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k})))\leq\frac{L(f(x^{0})-f^{\star})}{\bar{L}(1-\beta)(K+1)}.$ (19)
@@
$1+\bar{L}^{-2}\|\nabla f(x^{k^{\star}})\|^{2}\leq\left(\frac{L(f(x^{0})-f^{\star})}{\bar{L}(1-\beta)(K+1)}\right)^{2}+\frac{2L(f(x^{0})-f^{\star})}{\bar{L}(1-\beta)(K+1)}+1,$
@@
$\|\nabla f(x^{k^{\star}})\|^{2}\leq\left(\frac{L(f(x^{0})-f^{\star})}{(1-\beta)(K+1)}\right)^{2}+\frac{2\bar{L}L(f(x^{0})-f^{\star})}{(1-\beta)(K+1)}.$
Taking the square root and the fact that $\sqrt{\alpha+\beta}\leq\sqrt{\alpha}+\sqrt{\beta}$ we obtain the claimed result. ∎
@@
where we have used the fact that $\nabla f(x^{\star})=0$, $\nabla\phi^{*}(0)=0$, $\phi^{*}(0)=0$ and $\partial f^{*}(\nabla f(x))=\{x\}$. Taking once again the convex gradient inequality between points $\nabla f(x^{\star})$ and $\nabla f(x)$, we now have:
@@
$=-\|L^{-1}\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k}))\|^{2},$
where in the inequality we used Proposition 3.5 and in the equalities the fact that $\nabla\phi^{*}(L^{-1}\nabla f(x^{k}))=\frac{{h^{*}}^{\prime}(\|L^{-1}\nabla f(x^{k})\|)}{\|L^{-1}\nabla f(x^{k})\|}\bar{L}^{-1}\nabla f(x^{k})$. Plugging now this inequality into (22), we obtain the claimed result, which leads to
$\|x^{k+1}-x^{\star}\|^{2}\leq\|x^{k}-x^{\star}\|^{2}-L^{-1}\langle\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k})),x^{k}-x^{\star}\rangle\leq\|x^{k}-x^{*}\|^{2}-\|L^{-1}\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k}))\|^{2}$ (23)
Since $(\bar{L}h)^{*}=\bar{L}\star h^{*}$ and $h$ is an even function, we have that $(\bar{L}\phi)^{*}=\bar{L}\star(h^{*}\circ\|\cdot\|)$. In light of Proposition 2.2, we have that $f^{*}-L^{-1}(\bar{L}\phi)^{*}$ is a convex function. Then, since $f\in\mathcal{C}^{1}(\mathbb{R}^{n})$ and thus essentially smooth, from [maddison2021dual, Lemma 3.2] we have that [maddison2021dual, Assumption 3.1] holds. Therefore, substituting from [maddison2021dual, Equation (41)], we obtain $(\bar{L}\phi)^{*}(\nabla f(x^{k+1}))\leq(\bar{L}\phi)^{*}(\nabla f(x^{k}))$ for all $k\in\mathbb{N}_{0}$ and since $h^{*}$ is an increasing function on $\mathbb{R}_{+}$,
@@
This is the claimed result, since the function values decrease along the iterates of the algorithm. ∎
@@
To begin with, since $\operatorname{dom}\phi=\mathbb{R}^{n}$, and $f$ is $(L,\bar{L})$-anisotropically smooth, $f$ satisfies [laude2022anisotropic, Definition 3.1] relative to $\bar{L}\phi$ with constant $L$. Therefore, in light of [laude2022anisotropic, Proposition 4.1], $f=\inf_{y\in\mathbb{R}^{n}}\bar{L}(L^{-1}\star\phi)(\cdot-y)+\xi(y)$ for some $\xi:\mathbb{R}^{n}\to\overline{\mathbb{R}}$. Since $f$ is moreover convex in this setting, we can take $\xi$ to be convex and lsc, as described in the second part of [laude2022anisotropic, Proposition 4.1]. In order to prove our result, we will resort to a nonlinear proximal point interpretation of (2) with a strongly convex prox-kernel. We therefore consider the following iteration:
$x^{k+1}=\operatorname*{arg\,min}_{y}\bar{L}(L^{-1}\star\phi)(x^{k}-y)+\xi(y).$ (25)
From [laude2022anisotropic, Proposition 3.9 (ii)], we have that
@@
which directly implies that $x^{k+1}=x^{k}-L^{-1}\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k}))$ and the claimed equivalence between the two schemes is established. Using this result we can now prove a certain three-point-like property for the iterates generated by (2):
@@
Now, we move on to the proof of our main theorem. It is inspired by the proof of [doikov2020inexact], where we have also utilized the fact that $\phi$ is $2$-subhomogeneous.
@@
As established above, we consider the sequence of iterates generated by (25). In light of Lemma C.1 we have for any $x\in\mathbb{R}^{n}$:
@@
We choose $A_{k}=k^{2}$ and by using the fact that $\sum_{k=1}^{K}\tfrac{a_{k}^{2}}{A_{k}}\leq 4K$ [doikov2020inexact, Equation (35)]:
@@
which is known to have at most three solutions: $y_{1}<0$, $y_{2}=0$ and $y_{3}>0$. Without loss of generality we assume that $x>0$ and thus we get that there exists only one $\theta>0$ such that $g^{\prime}(\theta)=0$, which is a contradiction. ∎

### v2

@@ Appendix B Missing proofs of Section 2
We will prove that $T_{{L^{-1}},{\bar{L}^{-1}}}=\operatorname{id}-L^{-1}\nabla\phi^{*}\circ(\bar{L}^{-1}\nabla f)$ is a global homeomorphism from $\mathbb{R}^{n}\to\mathbb{R}^{n}$ using (Facchinei & Pang, 2003, Theorem 2.1.10).
Note that from Definition 2.4, $\lim_{\|x\|\to\infty}\|T_{{L^{-1}},{\bar{L}^{-1}}}(x)\|=\infty$, i.e. $T_{{L^{-1}},{\bar{L}^{-1}}}$ is norm-coercive, we only need to show that it is everywhere a local homeomorphism. The mapping $T_{{L^{-1}},{\bar{L}^{-1}}}$ is locally Lipschitz, since $\nabla\phi^{*}$ is globally Lipschitz and $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$ and thus the generalized Jacobian is well-defined. Now, we have that $\partial_{C}T_{{L^{-1}},{\bar{L}^{-1}}}(x)=\{I-L^{-1}V:V\in\partial_{C}(\nabla\phi^{*}\ \circ\bar{L}^{-1}\nabla f)(x)\}$ and in light of (Clarke, 1990, p. 75),
$\partial_{C}(\nabla\phi^{*}\circ\bar{L}^{-1}\nabla f)(x)v\subseteq\operatorname{con}\{\partial_{C}(\nabla\phi^{*})(\bar{L}^{-1}\nabla f(x))\bar{L}^{-1}\nabla^{2}f(x)v\}$
for any $v\in\mathbb{R}^{n}$. In order to show that $T_{{L^{-1}},{\bar{L}^{-1}}}$ is everywhere a local homeomorphism we are going to use Clarke’s inverse function theorem as presented in (Dontchev & Rockafellar, 2009, Theorem 4D.4). Consider thus any point $\bar{x}\in\mathbb{R}^{n}$ and the mapping $G_{A}(x)\coloneqq T_{{L^{-1}},{\bar{L}^{-1}}}(\bar{x})+A(x-\bar{x})$, where $A\in\partial_{C}T_{{L^{-1}},{\bar{L}^{-1}}}(\bar{x})$. Now, from the reasoning above we have that
$A(x-\bar{x})\in\operatorname{con}\{(I-L^{-1}\partial_{C}(\nabla\phi^{*})(\bar{L}^{-1}\nabla f(\bar{x}))\bar{L}^{-1}\nabla^{2}f(\bar{x}))(x-\bar{x})\}.$
From Definition 2.4 we have that for all $H\in\partial_{C}(\nabla\phi^{*})(\bar{L}^{-1}\nabla f(x))$, $\lambda_{\min}(I-L^{-1}\bar{L}^{-1}H\nabla^{2}f(\bar{x}))>0$, implying that $G_{A}$ is an invertible linear mapping for any $A\in\partial_{C}T_{{L^{-1}},{\bar{L}^{-1}}}(\bar{x})$. Therefore, from (Dontchev & Rockafellar, 2009, Theorem 4D.4), there exists a Lipschitz continuous mapping $T_{{L^{-1}},{\bar{L}^{-1}}}^{-1}$ such that $T_{{L^{-1}},{\bar{L}^{-1}}}^{-1}(T_{{L^{-1}},{\bar{L}^{-1}}}(x))=x$ for some neighborhoods $U$ of $\bar{x}$ and $V$ of $T_{{L^{-1}},{\bar{L}^{-1}}}(\bar{x})$. Since, moreover $T_{{L^{-1}},{\bar{L}^{-1}}}$ is Lipschitz continuous, it is a local homeomorphism at $\bar{x}$ in the sense of (Facchinei & Pang, 2003, Definition 2.1.9). Since this holds for any $\bar{x}\in\mathbb{R}^{n}$, it is everywhere a local homeomorphism. This concludes our proof by using (Facchinei & Pang, 2003, Theorem 2.1.10).
B.5 Proof of Corollary 2.7
Consider $\phi(x)=-\|x\|-\ln(1-\|x\|)$. In this case, $h(x)=-|x|-\ln(1-|x|)$ and thus from Lemma 1.3, $\nabla\phi^{*}(y)=\tfrac{y}{1+\|y\|}$. We thus have
$\nabla^{2}\phi^{*}(y)$ $=\frac{1}{1+\|y\|}I+\left(\frac{1}{(1+\|y\|)^{2}}-\frac{1}{1+\|y\|}\right)\frac{yy^{\top}}{\|y\|^{2}}$
$=\frac{1}{1+\|y\|}\left[I-\frac{\|y\|}{(1+\|y\|)}\frac{yy^{\top}}{\|y\|^{2}}\right]$
The term multiplying $\frac{yy^{\top}}{\|y\|^{2}}$ is negative and as such $\lambda_{\rm\max}(\nabla^{2}\phi^{*}(y))\leq\frac{1}{1+\|y\|}$. Moreover, $\nabla^{2}\phi^{*}(y)$ is positive-definite, since $1>\tfrac{\|y\|}{1+\|y\|}$, which then implies $\|\nabla^{2}\phi^{*}(y)\|\leq\frac{1}{1+\|y\|}$. Therefore,
$\|\nabla^{2}\phi^{*}(\bar{L}^{-1}y)\|\leq\frac{\bar{L}}{\bar{L}+\|y\|}=\frac{L_{0}}{L_{0}+L_{1}\|y\|},$
by choosing $\bar{L}=L_{0}/L_{1}$. By $(L_{0},L_{1})$-smoothness we moreover have $\|\nabla^{2}f(x)\|\leq L_{0}+L_{1}\|\nabla f(x)\|$ and thus
$\|\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))\nabla^{2}f(x)\|\leq\|\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))\|\|\nabla^{2}f(x)\|\leq L_{0}.$
Choosing now $L=L_{1}$ we further have
$\|\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))\nabla^{2}f(x)\|\leq L\bar{L}.$ (15)
Now take any points $x,\bar{x}\in\mathbb{R}^{n}$ and note that from the Cauchy–Schwarz inequality:
$\langle\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x))-\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\bar{x})),x-\bar{x}\rangle\leq\|\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x))-\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\bar{x}))\|\|x-\bar{x}\|.$ (16)
By the fundamental theorem of calculus for the mapping $\nabla\phi^{*}\circ(\bar{L}^{-1}\nabla f)$,
[…]
and as such
[…]
$\leq L\|x-\bar{x}\|,$
where the second inequality follows by (15). Putting this result back into (16), multiplying with $-\gamma<0$ and adding $\|x-\bar{x}\|^{2}$ to both sides we obtain:
@@
$\langle T_{{\delta L^{-1}},{\bar{L}^{-1}}}(x)-T_{{\delta L^{-1}},{\bar{L}^{-1}}}(\bar{x}),x-\bar{x}\rangle\geq(1-\delta)\|x-\bar{x}\|^{2},$
since $\gamma=\delta L^{-1}$.
@@
Note that (4) is equivalent to $\bar{x}\in\operatorname*{arg\,min}_{x\in\mathbb{R}^{n}}g(x):=\bar{L}(L^{-1}\star\phi)(x-\bar{y})-f(x)$ for all $\bar{x}\in\mathbb{R}^{n}$, where $\bar{y}=\bar{x}-L^{-1}\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\bar{x}))$. In light of the modern version of Fermat’s theorem (Rockafellar & Wets, 1998, Theorem 10.1), for $\tilde{x}$ to be a local minimizer of $g$, the following inclusion should hold: $0\in\widehat{\partial}g(\tilde{x})$. Through (Rockafellar & Wets, 1998, Exercise 10.10) this implies that
@@
or that $L(\tilde{x}-\bar{y})=\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\tilde{x}))$ or $\tilde{x}-L^{-1}\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\tilde{x}))=\bar{y}$. This means that $T_{{L^{-1}},{\bar{L}^{-1}}}(\tilde{x})=T_{{L^{-1}},{\bar{L}^{-1}}}(\bar{x})$ and since $T_{{L^{-1}},{\bar{L}^{-1}}}$ is injective, the only possible minimizer of $g$ is $\bar{x}$.
@@
Case 1: $\operatorname{dom}\phi$ is bounded. In light of (Rockafellar & Wets, 1998, p. 91), $g$ is a coercive function and as it is also proper and lsc, it attains its minimum.
@@
$\leq\bar{L}r^{-1}\alpha\phi\left(\frac{r}{\alpha}(x-\bar{y})\right)+\bar{L}r^{-1}(1-\alpha)\phi\left(\frac{r}{1-\alpha}\bar{y}\right)-\frac{\mu}{2}\bar{L}r^{-1}\alpha(1-\alpha)\left\|\frac{r}{\alpha}(x-\bar{y})-\frac{r}{1-\alpha}\bar{y}\right\|^{2},$
@@
Note now that $\psi$ is a proper, lsc and strongly convex function and as such it has bounded level-sets. Due to the inequality, the level-sets of $g$ are contained in those of $\psi$ and thus $g$ also has bounded level-sets. Since moreover $g$ is lsc, it attains its minimum.
@@ Appendix C Missing proofs of Section 3
$f(x^{k+1})\leq f(x^{k})-(1-|1-\alpha|)\bar{L}L^{-1}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k}))).$ (18)
@@
$\sum_{k=0}^{K}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k})))\leq\frac{L}{\bar{L}\beta}(f(x^{0})-f(x^{K+1}))\leq\frac{L}{\bar{L}\beta}(f(x^{0})-f^{\star}),$
@@
$(K+1)\min_{0\leq k\leq K}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k})))\leq\frac{L}{\bar{L}\beta}(f(x^{0})-f^{\star}).$ (19)
@@
$\min_{0\leq k\leq K}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k})))\leq\frac{L(f(x^{0})-f^{\star})}{\bar{L}\beta(K+1)}.$ (20)
@@
$1+\bar{L}^{-2}\|\nabla f(x^{k^{\star}})\|^{2}\leq\left(\frac{L(f(x^{0})-f^{\star})}{\bar{L}\beta(K+1)}\right)^{2}+\frac{2L(f(x^{0})-f^{\star})}{\bar{L}\beta(K+1)}+1,$
@@
$\|\nabla f(x^{k^{\star}})\|^{2}\leq\left(\frac{L(f(x^{0})-f^{\star})}{\beta(K+1)}\right)^{2}+\frac{2\bar{L}L(f(x^{0})-f^{\star})}{\beta(K+1)}.$
Taking the square root and using the fact that $\sqrt{\alpha+\beta}\leq\sqrt{\alpha}+\sqrt{\beta}$ we obtain the claimed result. ∎
@@
where we have used the fact that $\nabla f(x^{\star})=0$, $\nabla\phi^{*}(0)=0$, $\phi^{*}(0)=0$ and $x\in\partial f^{*}(\nabla f(x))$, since $f$ is convex. Taking once again the convex gradient inequality between points $\nabla f(x^{\star})$ and $\nabla f(x)$, we now have:
@@
$=-\|L^{-1}\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k}))\|^{2},$ (24)
where in the inequality we used Proposition 3.5 and in the equalities the fact that $\nabla\phi^{*}(L^{-1}\nabla f(x^{k}))=\frac{{h^{*}}^{\prime}(\|L^{-1}\nabla f(x^{k})\|)}{\|L^{-1}\nabla f(x^{k})\|}\bar{L}^{-1}\nabla f(x^{k})$ from Lemma 1.3. In the inequality, we also used the fact that ${h^{*}}^{\prime}(t)\geq 0$ for $t\geq 0$. Indeed, by convexity, we have that:
$h^{*}(0)\geq h^{*}(t)-{h^{*}}^{\prime}(t)t\Longleftrightarrow{h^{*}}^{\prime}(t)t\geq h^{*}(t)-h^{*}(0),$
implying that ${h^{*}}^{\prime}(t)\geq 0$ for all $t\geq 0$. Plugging now (C.4) into (23), we obtain
$\|x^{k+1}-x^{\star}\|^{2}\leq\|x^{k}-x^{\star}\|^{2}-L^{-1}\langle\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k})),x^{k}-x^{\star}\rangle\leq\|x^{k}-x^{*}\|^{2}-\|L^{-1}\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k}))\|^{2},$ (25)
which shows the Fejér monotonicity of $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ w.r.t. $x^{\star}\in\operatorname*{arg\,min}f$. Since $(\bar{L}h)^{*}=\bar{L}\star h^{*}$ and $h$ is an even function, we have that $(\bar{L}\phi)^{*}=\bar{L}\star(h^{*}\circ\|\cdot\|)$. In light of Proposition 2.2, we have that $f^{*}-L^{-1}(\bar{L}\phi)^{*}$ is a convex function. Now, for any $x,\bar{x}\in\mathbb{R}^{n}$, from the convex subgradient inequality for this function, between points $\nabla f(x)\in\operatorname{dom}\partial f^{*}$ and $\nabla f(\bar{x})\in\operatorname{dom}\partial f^{*}$ we have:
$(f^{*}-L^{-1}(\bar{L}\phi)^{*})(\nabla f(x))\geq(f^{*}-L^{-1}(\bar{L}\phi)^{*})(\nabla f(\bar{x}))+\langle\bar{x}-\nabla(L^{-1}(\bar{L}\phi)^{*})(\nabla f(\bar{x})),\nabla f(x)-\nabla f(\bar{x})\rangle,$
where we have moreover used the fact that $\bar{x}\in\partial f^{*}(\nabla f(\bar{x}))$. Therefore,
$D_{(\bar{L}\phi)^{*}}(\nabla f(x),\nabla f(\bar{x}))$ $=(\bar{L}\phi)^{*}(\nabla f(x))-(\bar{L}\phi)^{*}(\nabla f(\bar{x}))-\langle\nabla(\bar{L}\phi)^{*}(\nabla f(\bar{x})),\nabla f(x)-\nabla f(\bar{x})\rangle$
$\leq L[f^{*}(\nabla f(x))-f^{*}(\nabla f(\bar{x}))-\langle\bar{x},\nabla f(x)-\nabla f(\bar{x})\rangle]$
$=LD_{f}(\bar{x},x)$
where $D_{g}$ denotes the Bregman divergence associated with $g$ and the equality follows by the definition of the convex conjugate.
Thus, (Maddison et al., 2021, Assumption 3.1) holds for $k=(\bar{L}\phi)^{*}$. Therefore, substituting from (Maddison et al., 2021, Equation (41)), we obtain $(\bar{L}\phi)^{*}(\nabla f(x^{k+1}))\leq(\bar{L}\phi)^{*}(\nabla f(x^{k}))$ for all $k\in\mathbb{N}_{0}$ and since $h^{*}$ is an increasing function on $\mathbb{R}_{+}$ from Lemma 1.3,
@@
since $\frac{{h^{*}}^{\prime}(x)}{x}$ is decreasing on $\mathbb{R}_{+}$, implies that
@@
This is the claimed result, since the function values decrease along the iterates of the algorithm from (18). ∎
@@
In light of Proposition 2.2, $f=\inf_{y\in\mathbb{R}^{n}}\bar{L}(L^{-1}\star\phi)(\cdot-y)+\xi(y)$ for some $\xi:\mathbb{R}^{n}\to\overline{\mathbb{R}}$. Since $f$ is moreover convex in this setting and $\operatorname{dom}\phi=\mathbb{R}^{n}$, we can take $\xi$ to be convex and lsc from (Laude & Patrinos, 2025, Proposition 4.1). In order to prove our result, we will resort to a nonlinear proximal point interpretation of (2) with a strongly convex prox-kernel. We therefore consider the following iteration:
$x^{k+1}=\operatorname*{arg\,min}_{y\in\mathbb{R}^{n}}\bar{L}(L^{-1}\star\phi)(x^{k}-y)+\xi(y).$ (27)
From (Laude & Patrinos, 2025, Proposition 3.9 (ii)), since $\operatorname{dom}\phi=\mathbb{R}^{n}$, we have that
@@
which directly implies that $x^{k+1}=x^{k}-L^{-1}\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k}))$ and the claimed equivalence between the two schemes is established. Using this result we can now prove a certain three-point-like property for the iterates generated by (2).
@@
Now, we move on to the proof of our main theorem. It is inspired by the proof of (Doikov & Nesterov, 2020), where we have also utilized the fact that $\phi$ is $2$-subhomogeneous.
@@
As established above, we consider the sequence of iterates generated by (27). In light of Lemma C.1 and since $\phi\geq 0$ we have for any $x\in\mathbb{R}^{n}$:
@@
We choose $A_{k}=k^{2}$ and by using the fact that $\sum_{k=1}^{K}\tfrac{a_{k}^{2}}{A_{k}}\leq 4K$ (Doikov & Nesterov, 2020, Equation (35)):
@@
which has exactly three solutions: $y_{1}<0$, $y_{2}=0$, $y_{3}>0$, since $2\cosh(x)>2+x^{2}$ for $x\neq 0$. Without loss of generality we assume that $x>0$ and thus we get that there exists only one $\theta>0$ such that $g^{\prime}(\theta)=0$, which is a contradiction. ∎